OpenStudy (anonymous):

13. 5/17n=-9 (1 point)

OpenStudy (anonymous):

n = -153/5

OpenStudy (anonymous):

can u explain please mate

OpenStudy (anonymous):

(5/17)n = -9 5n = -9*17 5n = -153 n = -153/5